OpenStudy (anonymous):

Your piggy bank has a total of 43 coins in it; some are dimes and some are nickels. If you have a total of $3.60, how many dimes do you have? Numerical Answers Expected!

OpenStudy (tennistar):

Let d = dimes, n = nickles, and money in terms of cents # dimes + # nickels = 43 d + n = 43 dime total + nickel total = 360 10d + 5n = 360 You end up with two equations: 10d + 5n = 360 d + n = 43 What values of d and n will make the above true? 10d + 5n = 360 10d + 10n = 430 {multiplied both sides by 10} -5n = -70 {subtracted bottom equation from top equation} n = 14 {There are 14 nickels} d + n = 43 d + 14 = 43 d = 29 {There are 29 dimes}

OpenStudy (tennistar):

Are you taking FLVS ALGebra

OpenStudy (anonymous):

Why?

OpenStudy (tennistar):

Well, I am too

OpenStudy (tennistar):

I have seen these questions before

OpenStudy (anonymous):

Yeah i was taking it but now im doing a worksheet

OpenStudy (tennistar):

oh ok